OpenStudy (anonymous):

Integrating two ways, one with a u sub only and one with a different u sub and then by parts I am out by more than just a constant. Integrand is x^5(x^3+1)^1/2 so first u sub was u=x^3 +1 and in second attempt u=x^3...help

OpenStudy (anonymous):

A somewhat tricky substitution: let \(u^{2/3}=x~\iff~u^2=x^3\), so that \(dx=\dfrac{2}{3}u^{-1/3}~du\). \[\int x^5\sqrt{x^3+1}~dx~~\iff~~\int u^{10/3}\sqrt{u^2+1}\left(\frac{2}{3}u^{-1/3}\right)~du\] Simplifying a bit leaves you with \[\frac{2}{3}\int u^{9/3}\sqrt{u^2+1}~du=\frac{2}{3}\int u^3\sqrt{u^2+1}~du\] which looks like it's easier to work with. You might have to find a similar but different substitution like the one above.

OpenStudy (anonymous):

Or perhaps a trigonometric substitution will help right away: \(u=\tan t\), so you have \(du=\sec^2t~dt\) and \[\frac{2}{3}\int\tan^3t\sqrt{\tan^2t+1}~\sec^2t~dt=\frac{2}{3}\int\tan^3t\sec^3t~dt\] And this I'm sure can be done, with some effort.

OpenStudy (anonymous):

I made a chrome extension that makes it easy to copy what other people have posted: \(\color{blue}{\text{Originally Posted by}}\) @SithsAndGiggles A somewhat tricky substitution: let \(u^{2/3}=x~\iff~u^2=x^3\), so that \(dx=\dfrac{2}{3}u^{-1/3}~du\). \[\int x^5\sqrt{x^3+1}~dx~~\iff~~\int u^{10/3}\sqrt{u^2+1}\left(\frac{2}{3}u^{-1/3}\right)~du\] Simplifying a bit leaves you with \[\frac{2}{3}\int u^{9/3}\sqrt{u^2+1}~du=\frac{2}{3}\int u^3\sqrt{u^2+1}~du\] which looks like it's easier to work with. You might have to find a similar but different substitution like the one above. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

@wio, neat! I'm surprised OS doesn't already have a feature like this

OpenStudy (anonymous):

Yeah, it was something that I wish existed because for my more latex heavy posts it's hard to retype and correct them.

OpenStudy (anonymous):

Sorry for invading @catanary. I hope you solve the problem.

OpenStudy (anonymous):

I think we've got in under control. Thanks!

OpenStudy (anonymous):

The substitution u=x^3 +1 then du = 3x^2 and x^3=u-1 gives 1/3∫(u-1)√u du which gives the correct answer \[\frac{ -2 }{ 9 }\left( x^{3} +1 \right)^{\frac{ 3 }{2 }}+\frac{ 2 }{ 15 }\left( x^{3} +1 \right)^{\frac{ 5 }{ 2 }}\] but I am not sure why when I let u = x^3 then du is 3x^2 so integral is 1/3∫u√u+1du and do this by parts I get an x^3 in my asnwer which is not supposed to be there.

OpenStudy (anonymous):

Are you getting something like \[\frac{2}{3}x^3\left(x^3+1\right)^{3/2}-\frac{4}{15}\left(x^3+1\right)^{5/2}+C~~?\]

OpenStudy (anonymous):

That' exactly what I am getting...Please can you tell me why its wrong?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

Hmm, according to Wolfram, the by-parts solution is off by only a factor of 3: http://www.wolframalpha.com/input/?i=D%5B%282%2F3%29*x%5E3*%28x%5E3%2B1%29%5E%283%2F2%29-%284%2F15%29%28x%5E3%2B1%29%5E%285%2F2%29%2Cx%5D I would say there's an algebraic error in your work and in mine, but I don't see where it is...